# On stable assignments generated by choice functions of mixed type

### v1

@@ 6 The poset of rotations
Summing up the above observations, we can conclude that the semi-route $X$ can be transformed so that the rotation $\rho$ reaches the first position in the resulting sequence of rotations related to $z$ and $x$. Transforming $Y$ in a similar way, we obtain a sequence of rotations related to $z$ and $y$ in which the same rotation $\rho$ is situated at the first position as well. Then we act as in the beginning of the proof, yielding a contradiction with the definition of $z$. This proves ((6.2)), implying the left relation in (6.1). The right relation in (6.1) is proved symmetrically (by taking the order $\prec_{F}$ instead of $\prec_{W}$, and thinking of a route as a sequence going from $x^{\rm max}$ to $x^{\rm min}$). Thus, $\omega$ gives a homomorphism of the above lattices as required.
Now to see that $\mathfrak{C}$ is bijective to ${\cal S}$, it suffices to note that for any two rotations $\rho$ and $\rho^{\prime}$, there is $x\in{\cal S}$ which “separates” these rotations, in the sense that a lower semi-route for $x$ uses one of $\rho,\rho^{\prime}$, while an upper semi-route for $x$ uses the other of them. $\Box$
@@
For $v\in V$, let $A_{v}$ and $A^{\prime}_{v}$ denote the sets of active edges leaving $v$ in $\overrightarrow{\Gamma}$ and $\overrightarrow{\Gamma}^{\prime}$, respectively. Then (cf. ((5.2))(b)):
@@
    if $v=f\in F$, then either $\pi^{\rm c}_{f}(\widetilde{x}^{\prime})=\pi^{\rm c}_{f}(\widetilde{x})$ and $A^{\prime}_{f}\subseteq A_{f}$, or $\pi^{\rm c}_{f}(\widetilde{x}^{\prime})$ is worse than $\pi^{\rm c}_{f}(\widetilde{x})$; whereas if $v=w\in W$, then either $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})=\pi^{\rm c}_{w}(\widetilde{x})$ and $A^{\prime}_{w}\supseteq A_{w}$, or $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})$ is better than $\pi^{\rm c}_{w}(\widetilde{x})$; moreover, if $\pi^{\rm c}_{v}(\widetilde{x}^{\prime})=\pi^{\rm c}_{v}(\widetilde{x})$ for all $v\in V$, then at least one of the above inclusions is strict.
@@
Case 1: $L$ and $L^{\prime}$ have a common vertex. For each $v\in V_{L}\cap V_{L^{\prime}}$, we have: $A_{v}=E_{L}\cap E_{v}$ and $A^{\prime}_{v}=E_{L^{\prime}}\cap E_{v}$, the sets $A_{v},\,A^{\prime}_{v},\,\pi^{\rm c}_{v}(\widetilde{x}),\,\pi^{\rm c}_{v}(\widetilde{x}^{\prime})$ satisfy relations as in ((6.3)), and at least one of these relations is different from equality (since, obviously, $L\neq L^{\prime}$). When moving along any route, relations concerning the active graphs for neighboring assignments are updated similar to those exposed in ((6.3)). These updates have a “monotone character” at each vertex, and we can conclude that the rotation $\rho^{\prime}$ can never be applied in a route earlier than $\rho$. It follows that $\rho\lessdot\rho^{\prime}$.
@@
Case 2: $L$ and $L^{\prime}$ are disjoint. This is possible only if $L^{\prime}$ is a strong component in each of $\overrightarrow{\Gamma}$ and $\overrightarrow{\Gamma}^{\prime}$, but $L^{\prime}$ is maximal in $\overrightarrow{\Gamma}^{\prime}$, and not maximal in $\overrightarrow{\Gamma}$. Then the set $Q$ of active edges in $\overrightarrow{\Gamma}$ going from $L^{\prime}$ to other strong components in $\overrightarrow{\Gamma}$ must disappear under the transformation of $\widetilde{x}$ along $\rho$. This implies that all edges in $Q$ enter just $L$. Since the assignment on each $e\in Q$ is not changed: $\widetilde{x}^{\prime}(e)=\widetilde{x}(e)$, the situation that the edge $e$ becomes non-active can happen only if $e$ goes from $F$ to $W$, say, $e=(f,w)$, and $w$ does not belong to the tail $T_{w}(\widetilde{x}^{\prime})$. This implies that the assignments on the edges in $\pi^{\rm c}_{w}(\widetilde{x})$ must reduce to zero; then the critical tie in $\Pi_{w}$ becomes better: $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})>\pi^{\rm c}_{w}(\widetilde{x})$. Moreover, $e$ must belong to a tie worse than $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})$. It follows that in any route the rotation $\rho^{\prime}$ cannot be applied earlier than $\rho$ (for otherwise the edges from $Q$ should be active at the moment of applying $\rho^{\prime}$, and therefore, $L^{\prime}$ could not be a maximal strong component in the current active graph). As a result, we obtain $\rho\lessdot\rho^{\prime}$. $\Box$
@@
In the rest of this section we explain how to construct the rotation poset $({\cal R},\tau,\lessdot)$ efficiently. Our method is of a similar spirit as that elaborated for the stable allocation problem in [4] (and earlier for the stable marriage problem in [9]).
@@
First of all we compute, step by step, the set ${\cal R}$ of rotations, see explanations around Proposition 5.1. This is performed in parallel with the construction of a route $T$ from $x^{\rm min}$ to $x^{\rm max}$. Recall that at each step, for a current stable assignment $x$, we construct the active graph $\overrightarrow{\Gamma}(x)$, extract the set ${\cal L}^{\rm rot}(x)$ of maximal strong components $L$ in it, and determine the rotations $\rho=\rho_{L}$ and their maximal admissible weights $\tau(\rho)=\tau_{L}$. (Here finding ${\cal L}^{\rm rot}(x)$ can be performed in linear time $O(|E|)$, while computing the rotations $\rho_{L}$, which is a more involved task, takes time polynomial in $|E|$.) Then we update the current $x$ by adding to it the total shift $\sum(\tau_{L}\rho_{L}\colon L\in{\cal L}^{\rm rot}(x))$, and so on. As a result, we obtain the set ${\cal R}$ of all rotations $\rho$ (where $|{\cal R}|=O(|E|)$), along with their maximal weights $\tau(\rho)$, in time polynomial in $|E|$.
@@
    there exists $x\in{\cal S}$ such that $\rho$ but not $\rho^{\prime}$ is applicable to $x$, whereas $\rho^{\prime}$ is applicable to the assignment $x^{\prime}:=x+\tau(\rho)\rho$.
One can see that each pair $(\rho,\rho^{\prime})$ in ${\cal E}$ gives an immediately preceding relation. (Indeed, $x$ and $x^{\prime}$ as above can be represented as two consecutive elements in some route; also the Claim implies that $\rho\lessdot\rho^{\prime}$; therefore, the existence of $\rho^{\prime\prime}$ between $\rho$ and $\rho^{\prime}$ is impossible.)
@@
We show that ${\cal E}$ does determine all immediately preceding relations in $({\cal R},\lessdot)$ (and therefore, the directed graph $({\cal R},{\cal E})$ can be used to encode this poset and, moreover, to store the information about all stable assignments). Clearly if $\rho^{\prime}\in{\cal R}$ is reachable from $\rho\in{\cal R}$ by a directed path in $({\cal R},{\cal E})$, then $\rho\lessdot\rho^{\prime}$. The converse is valid as well.
@@
For $\rho,\rho^{\prime}\in{\cal R}$, if $\rho\lessdot\rho^{\prime}$, then the graph ${\cal H}=({\cal R},{\cal E})$ has a directed path from $\rho$ to $\rho^{\prime}$.
Proof Let ${\cal I}$ be the set of ideals in the poset $({\cal R},\lessdot)$, and ${\cal J}$ the set of ideals $I$ in the poset determined by ${\cal H}$ (via the reachability by paths). Then ${\cal I}\subseteq{\cal J}$, and we have to show that ${\cal I}={\cal J}$, whence the result will follow.
@@
Consider a rotation $\rho$ applicable to $x$. Then $R^{\prime}:=R\cup\{\rho\}$ is an ideal in ${\cal I}$ (corresponding to the assignment $x^{\prime}:=x+\tau(\rho)\rho$). By the maximality of $R$, $\rho$ is outside $J$. We have $J^{\prime}:=J\cup\{\rho\}\in{\cal J}$, and the maximality of $J$ implies $J^{\prime}\in{\cal I}$.
@@
Now since $J^{\prime}\supset R^{\prime}$ and both $J^{\prime},R^{\prime}$ are ideals in ${\cal I}$, the set $J^{\prime}-R^{\prime}$ contains a rotation $\rho^{\prime}$ such that the set $R^{\prime\prime}:=R^{\prime}\cup\{\rho^{\prime}\}$ is an ideal in ${\cal I}$. At the same time, $J^{\prime\prime}:=R\cup\{\rho^{\prime}\}$ is not an ideal in ${\cal I}$; for otherwise, in view of $R\subset J^{\prime\prime}\subseteq J$, we would obtain a contradiction with the maximality of $R$ (if $J^{\prime\prime}\neq J$) or with the condition on $J$ (if $J^{\prime\prime}=J$).
@@
So we come to the situation when $R,R^{\prime},R^{\prime\prime}\in{\cal I}$, $R^{\prime}=R\cup\{\rho\}$, $R^{\prime\prime}=R^{\prime}\cup\{\rho^{\prime}\}$, but $R\cup\{\rho^{\prime}\}\notin{\cal I}$. Then the rotation $\rho^{\prime}$ is not applicable to the assignment $x$, but it is applicable to $x^{\prime}$ (obtained from $x$ by applying $\rho$). This means that $(\rho,\rho^{\prime})$ is an edge in ${\cal E}$ (cf. ((6.4))). But this edge goes from ${\cal R}-J$ to $J$ (since $\rho\notin J$ while $\rho^{\prime}\in J$), contrary to the fact $J$ is an ideal for ${\cal H}$. $\Box$

### v2

@@ 6 The poset of rotations
Summing up the above observations, we can conclude that the semi-route $X$ can be transformed so that the rotation $\rho$ reaches the first position in the resulting sequence of rotations related to $z$ and $x$. By transforming $Y$ in a similar way, we obtain a sequence of rotations related to $z$ and $y$ in which the same rotation $\rho$ is situated at the first position as well. Then we act as in the beginning of the proof, yielding a contradiction with the definition of $z$. This proves ((6.2)), implying the left equality in (6.1). The right equality in (6.1) is proved symmetrically (by taking the order $\prec_{F}$ instead of $\prec_{W}$, and thinking of a route as a sequence going from $x^{\rm max}$ to $x^{\rm min}$). Thus, $\omega$ gives a homomorphism of the above lattices, as required.
It follows that the map $\omega$ homomorphically embeds $(S,\prec_{W})$ into a sublattice of $(\mathfrak{C},<)$. Also $\omega$ is, obviously, injective.
To show that $\omega$ is a bijection, we use the fact that for any $\rho,\rho^{\prime}\in{\cal R}$, unless $\rho^{\prime}\lessdot\rho$, there is a route where $\rho$ is applied earlier than $\rho^{\prime}$, and therefore, there is $x\in{\cal S}$ separating $\rho$ from $\rho^{\prime}$, in the sense that $\omega(x)$ assigns a positive weight to $\rho$, but zero weight to $\rho^{\prime}$. This implies (by using (6.1)(left)) that for any $\rho\in{\cal R}$ and $0<\lambda\leq\tau(\rho)$, there is $x\in{\cal S}$ such that $\omega(x)$ assigns weight $\lambda$ to $\rho$, the full weight $\tau(\rho^{\prime\prime})$ to each $\rho^{\prime\prime}\lessdot\rho$, and zero weight otherwise (in other words, $\omega(x)$ is the minimal closed function taking value $\lambda$ on $\rho$). So $\omega({\cal S})$ covers all join-irreducible elements of $(\mathfrak{C},<)$, and the desired bijection follows. $\Box$
Remark 4.  By Birkhoff’s representation theorem [4], a finite distributive lattice ${\cal L}$ is isomorphic to the lattice (“ring of sets”), with standard $\cap$ and $\cup$ operations, formed by ideals (down sets) in the poset ${\cal P}$ generated by join-irreducible elements of ${\cal L}$. This can be extended, with a due care, to our infinite lattice ${\cal L}=({\cal S},\prec_{W})$ as well. Formally, for an integer $k\geq 2$ and any $\rho\in{\cal R}$, choose a set $\Lambda^{(k)}(\rho)$ of $k$ weights $0=\lambda_{1}<\lambda_{2}<\cdots<\lambda_{k}=\tau(\rho)$. Accordingly, transform the weighted poset $({\cal R},\tau,\lessdot)$ into a finite poset ${\cal P}^{(k)}$ by replacing each element $\rho$ by $k$ elements $\rho_{1},\ldots,\rho_{k}$ corresponding to the weights in $\Lambda^{(k)}(\rho)$ and ordered as $\rho_{1}\lessdot\cdots\lessdot\rho_{k}$, respectively (as to relations $\rho\lessdot\rho^{\prime}$ in $({\cal R},\tau,\lessdot)$, they are replaced by $\rho_{k}\lessdot\rho^{\prime}_{1}$). Applying the rotational shifts with chosen weights, we build the corresponding finite subset ${\cal S}^{(k)}$ of stable assignments in ${\cal S}$. This ${\cal S}^{(k)}$ induces a sublattice of $({\cal S},\prec_{W})$ isomorphic to the lattice of ideals in ${\cal P}^{(k)}$, in spirit of Birkhoff’s representation. (Note that ${\cal S}^{(2)}$ consists exactly of the stable assignments generated by full rotational shifts, and ${\cal P}^{(2)}$ turns into the unweighted poset $({\cal R},\lessdot)$.) When $k$ grows, the constructed objects “tend” to the original objects ${\cal S},\lessdot_{W},{\cal R},\tau,\lessdot$, and one can conclude that the join-irreducible elements $x$ of $({\cal S},\prec_{W})$ are exactly the preimages by $\omega$ of the minimal (viz. principal) closed functions determined by the pairs $(\rho,\lambda)$ with $\rho\in{\cal R}$ and $0<\lambda\leq\tau(\rho)$.
@@
For $v\in V^{=}$, let $A_{v}$ and $A^{\prime}_{v}$ denote the sets of active edges leaving $v$ in $\overrightarrow{\Gamma}$ and $\overrightarrow{\Gamma}^{\prime}$, respectively. Then (cf. ((5.2))(b)):
@@
    if $v=f\in F^{=}$, then either $\pi^{\rm c}_{f}(\widetilde{x}^{\prime})=\pi^{\rm c}_{f}(\widetilde{x})$ and $A^{\prime}_{f}\subseteq A_{f}$, or $\pi^{\rm c}_{f}(\widetilde{x}^{\prime})$ is worse than $\pi^{\rm c}_{f}(\widetilde{x})$; whereas if $v=w\in W^{=}$, then either $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})=\pi^{\rm c}_{w}(\widetilde{x})$ and $A^{\prime}_{w}\supseteq A_{w}$, or $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})$ is better than $\pi^{\rm c}_{w}(\widetilde{x})$; moreover, if $\pi^{\rm c}_{v}(\widetilde{x}^{\prime})=\pi^{\rm c}_{v}(\widetilde{x})$ for all $v\in V^{=}$, then at least one of the above inclusions is strict.
@@
Case 1: $L$ and $L^{\prime}$ have a common vertex. For each $v\in V_{L}\cap V_{L^{\prime}}$, we have: $A_{v}=E_{L}\cap E_{v}$ and $A^{\prime}_{v}=E_{L^{\prime}}\cap E_{v}$, the sets $A_{v},\,A^{\prime}_{v},\,\pi^{\rm c}_{v}(\widetilde{x}),\,\pi^{\rm c}_{v}(\widetilde{x}^{\prime})$ satisfy relations as in ((6.3)), and at least one of these relations does not turn into equality (since, obviously, $L\neq L^{\prime}$). When moving along any route, relations concerning the active graphs for neighboring assignments are updated similar to those exposed in ((6.3)). These updates have a “monotone character” at each vertex, and we can conclude that the rotation $\rho^{\prime}$ cannot be applied in any route earlier than $\rho$. It follows that $\rho\lessdot\rho^{\prime}$.
@@
Case 2: $L$ and $L^{\prime}$ are disjoint. This is possible only if $L^{\prime}$ is a strong component in each of $\overrightarrow{\Gamma}$ and $\overrightarrow{\Gamma}^{\prime}$, but $L^{\prime}$ is maximal in $\overrightarrow{\Gamma}^{\prime}$ and not maximal in $\overrightarrow{\Gamma}$. Then the set $Q$ of active edges in $\overrightarrow{\Gamma}$ going from $L^{\prime}$ to other strong components in $\overrightarrow{\Gamma}$ must disappear under the transformation of $\widetilde{x}$ into $\widetilde{x}^{\prime}$. This implies that all edges in $Q$ enter just $L$. Since the assignment on each $e\in Q$ is not changed: $\widetilde{x}^{\prime}(e)=\widetilde{x}(e)$, the situation that the edge $e$ becomes non-active can happen only if $e$ goes from $F$ to $W$, say, $e=(f,w)$, and $w$ does not belong to the tail $T_{w}(\widetilde{x}^{\prime})$. This implies that the assignments on the edges in $\pi^{\rm c}_{w}(\widetilde{x})$ must reduce to zero. Then the critical tie in $\Pi_{w}$ becomes better: $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})>\pi^{\rm c}_{w}(\widetilde{x})$. Moreover, $e$ must belong to a tie worse than $\pi^{\rm c}_{w}(\widetilde{x}^{\prime})$. It follows that in any route the rotation $\rho^{\prime}$ cannot be applied earlier than $\rho$ (for otherwise the edges from $Q$ should be active at the moment of applying $\rho^{\prime}$, and therefore, $L^{\prime}$ could not be a maximal strong component in the current active graph). As a result, we obtain $\rho\lessdot\rho^{\prime}$. $\Box$
@@
In the rest of this section, we explain how to construct the rotation poset $({\cal R},\tau,\lessdot)$ efficiently. Our approach is in spirit of the method elaborated for the stable marriage problem in [12].
@@
First of all we compute, step by step, the set ${\cal R}$ of rotations, see explanations around Proposition 5.1. This is performed in the process of constructing an arbitrary route $T$ from $x^{\rm min}$ to $x^{\rm max}$. Recall that at each step, for a current stable assignment $x$, we construct the active graph $\overrightarrow{\Gamma}(x)$, extract the set ${\cal L}^{\rm rot}(x)$ of maximal strong components $L$ in it, and determine the rotations $\rho=\rho_{L}$ and their maximal admissible weights $\tau(\rho)=\tau_{L}$ (this is performed in time polynomial in $|E|$, cf. Proposition 5.1). Then we update the current $x$ by adding to it the total shift $\sum(\tau_{L}\rho_{L}\colon L\in{\cal L}^{\rm rot}(x))$, and so on. As a result, we obtain the set ${\cal R}$ of all rotations $\rho$ (where $|{\cal R}|=O(|E|)$) along with their maximal weights $\tau(\rho)$, in time polynomial in $|E|$.
@@
    there exists $x\in{\cal S}$ such that $\rho$ but not $\rho^{\prime}$ is applicable to $x$, whereas $\rho^{\prime}$ is applicable to $x^{\prime}:=x+\tau(\rho)\rho$.
Each pair $(\rho,\rho^{\prime})\in{\cal E}$ gives an immediately preceding relation. (Indeed, $x$ and $x^{\prime}$ as above can be represented as two consecutive elements in some route; also the Claim implies that $\rho\lessdot\rho^{\prime}$; therefore, the existence of $\rho^{\prime\prime}$ between $\rho$ and $\rho^{\prime}$ is impossible.)
@@
We show that ${\cal E}$ does determine all immediately preceding relations in $({\cal R},\lessdot)$ (in other words, $({\cal R},{\cal E})$ is the Hasse diagram for this poset). Clearly if $\rho^{\prime}\in{\cal R}$ is reachable from $\rho\in{\cal R}$ by a directed path in $({\cal R},{\cal E})$, then $\rho\lessdot\rho^{\prime}$. The converse is valid as well.
@@
If $\rho,\rho^{\prime}\in{\cal R}$ and $\rho\lessdot\rho^{\prime}$, then the graph ${\cal H}=({\cal R},{\cal E})$ has a directed path from $\rho$ to $\rho^{\prime}$.
Proof  Let ${\cal I}$ be the set of ideals in the poset $({\cal R},\lessdot)$, and ${\cal J}$ the set of ideals in the poset determined by ${\cal H}$ (via the reachability by paths). Then ${\cal I}\subseteq{\cal J}$, and we have to show that ${\cal I}={\cal J}$, whence the result will follow.
@@
Consider a rotation $\rho$ applicable to $x$. Then $R^{\prime}:=R\cup\{\rho\}$ is an ideal in ${\cal I}$ (corresponding to the assignment $x^{\prime}:=x+\tau(\rho)\rho$). By the maximality of $R$, we have $R^{\prime}\not\subseteq J$; so $\rho$ is not in $J$. Then the set $J^{\prime}:=J\cup R^{\prime}=J\cup\{\rho\}$ is an ideal in ${\cal J}$ strictly including $J$, and the maximality of $J$ implies $J^{\prime}\in{\cal I}$.
@@
Now since $J^{\prime}\supset R^{\prime}$ and both $J^{\prime},R^{\prime}$ are ideals in ${\cal I}$, the set $J^{\prime}-R^{\prime}$ contains a rotation $\rho^{\prime}$ such that the set $R^{\prime\prime}:=R^{\prime}\cup\{\rho^{\prime}\}$ is an ideal in ${\cal I}$. Form $J^{\prime\prime}:=R\cup\{\rho^{\prime}\}$; obviously, $R\subset J^{\prime\prime}\subseteq J$. This $J^{\prime\prime}$ is not an ideal in ${\cal I}$ (since $J^{\prime\prime}\in{\cal I}$ would lead to a contradiction with the maximality of $R$ (if $J^{\prime\prime}\neq J$) or with the condition on $J$ (if $J^{\prime\prime}=J$)).
@@
So we come to the situation when $R,R^{\prime},R^{\prime\prime}\in{\cal I}$; $R^{\prime}=R\cup\{\rho\}$; $R^{\prime\prime}=R^{\prime}\cup\{\rho^{\prime}\}$; but $R\cup\{\rho^{\prime}\}\notin{\cal I}$. Then the rotation $\rho^{\prime}$ is not applicable to the assignment $x$, but it is applicable to $x^{\prime}$ (obtained from $x$ by applying $\rho$). This means that $(\rho,\rho^{\prime})$ is an edge in ${\cal E}$ (cf. ((6.4))). But this edge goes from ${\cal R}-J$ to $J$ (since $\rho\notin J$ while $\rho^{\prime}\in J$), contrary to the fact that $J$ is an ideal for ${\cal H}$. $\Box$
7 Affine representability and stable assignments of minimum cost
The bijection $\omega$ figured in Proposition 6.3 enables us to give an affine representation of the lattice ${\cal L}=({\cal S},\prec_{W})$, like what is noticed in [18] for SAP, or in [7] for the Boolean case of Alkan–Gale’s model.
In what follows, depending on the context, a rotation $\rho\in{\cal R}$ can be regarded either as the corresponding function on $E$ (viz. a vector in ${\mathbb{R}}^{E}$), or as a unit base vector in the space ${\mathbb{R}}^{\cal R}$ (of which coordinates are indexed by elements of ${\cal R}$); in the latter case, we may use notation $\langle\rho\rangle$ for $\rho$. Let $A\in{\mathbb{R}}^{E\times{\cal R}}$ be the matrix whose columns are formed by rotations $\rho\in{\cal R}$ (regarded as vectors in ${\mathbb{R}}^{E}$). Then, by Proposition 6.3,
  1. (7.1)
    the affine map $\gamma:\lambda\in{\mathbb{R}}^{\cal R}\mapsto x\in{\mathbb{R}}^{E}$ given by $x:=x^{\rm min}+A\lambda$ determines the bijection between the lattice $(\mathfrak{C},<)$ of closed functions $\lambda$ for the rotation poset $({\cal R},\tau,\lessdot)$ and the lattice $({\cal S},\prec_{W})$ of stable assignments in SMP; in other words, $x=\omega^{-1}(\lambda)$ (where $x\in{\cal S}$, $\lambda\in\mathfrak{C}$, and $\omega$ is as in Proposition 6.3).
As is explained in Sects. 5 and 6, $|{\cal R}|$ is estimated as $O(|E|)$ and the matrix $A$ can be constructed in time polynomial in $|E|$. Due to this, we can take advantages from the above affine representation in order to efficiently translate certain problems on stable assignments into problems on closed functions.
First of all we observe that the map $\gamma$ in ((7.1)) sends the convex hull ${\rm conv}(\mathfrak{C})\subset{\mathbb{R}}^{\cal R}$ of $\mathfrak{C}$ onto the convex hull ${\rm conv}({\cal S})\subset{\mathbb{R}}^{E}$ of ${\cal S}$. The polytope ${\rm conv}(\mathfrak{C})$ is full-dimensional in ${\mathbb{R}}^{\cal R}$ (since each unit base vector $\langle\rho\rangle$ can be expressed as a linear combination of closed functions) and it can be described by the linear constraints
$0\leq\lambda(\rho)\leq\tau(\rho),\qquad\quad$ $\rho\in{\cal R};$ (7.2)
$(1/\tau(\rho))\lambda(\rho)\leq(1/\tau(\rho^{\prime}))\lambda(\rho^{\prime}),$ $\rho,\rho^{\prime}\in{\cal R},\;\rho^{\prime}\lessdot\rho.$ (7.3)
(Note that (7.2)–(7.3) turn into the constraints for Stanley’s order polytope [22] after scaling each variable $\lambda(\rho)$ to $\lambda(\rho)/\tau(\rho)$.)
When $A$ has full column rank (and therefore, ${\rm dim}({\rm conv}({\cal S}))={\rm dim}({\rm conv}(\mathfrak{C}))=|{\cal R}|$), ${\rm conv}({\cal S})$ becomes affinely congruent to ${\rm conv}(\mathfrak{C})$, and the facets of ${\rm conv}({\cal S})$ can be explicitly described via the facets of ${\rm conv}(\mathfrak{C})$; the latter ones are as in (7.2) for minimal $\rho$ in $({\cal R},\lessdot)$, and as in (7.3) when $\rho^{\prime}$ immediately precedes $\rho$. This is valid, e.g., for the Boolean case of [1], as is shown in [7], and for the unconstrained case of SAP (when $b\equiv\infty)$), as is shown in [18]. As to our model SMP, the full column rank is not guaranteed even in the unconstrained case $b\equiv\infty$ (to construct a counterexample is not difficult). When $k:=|{\cal R}|-{\rm dim}({\rm conv}({\cal S}))>0$, the facets of ${\rm conv}({\cal S})$ are images by $\gamma$ of some faces (or their unions) of co-dimension $k+1$ in ${\rm conv}(\mathfrak{C})$, and to characterize these facets looks a less trivial task (for a further discussion, see Remark 5 below).
Next, note that
  1. (7.4)
    the vertices of ${\rm conv}(\mathfrak{C})$ are exactly the fully closed functions.
Indeed, if $\lambda\in\mathfrak{C}$ is such that $0<\lambda(\rho)<\tau(\rho)$ holds for some $\rho\in{\cal R}$, then $\lambda$ is a convex combination of closed functions $\lambda^{\prime}$ and $\lambda^{\prime\prime}$ obtained from $\lambda$ by replacing its value on $\rho$ by zero and by $\tau(\rho)$, respectively; so $\lambda$ is not a vertex. Conversely, any fully closed function $\lambda$ is a vertex of ${\rm conv}(\mathfrak{C})$ since it is the unique element of $\mathfrak{C}$ attaining the maximum of the inner product $a\lambda^{\prime}$ over $\lambda^{\prime}\in\mathfrak{C}$, where for $\rho\in{\cal R}$, $a(\rho)$ is defined to be 1 if $\lambda(\rho)=\tau(\rho)$, and $-1$ if $\lambda(\rho)=0$.
We obtain from ((7.1)) and ((7.4)) that
  1. (7.5)
    if $x\in{\cal S}$ is a vertex of ${\rm conv}({\cal S})$, then $\omega(x)$ is fully closed (but the converse is not guaranteed when $\gamma$ is singular).
In the rest of this section we consider the min-cost stable assignment problem:
  1. (7.6)
    given costs $c(e)\in{\mathbb{R}}$ of edges $e\in E$, find a stable assignment $x\in{\cal S}$ minimizing the total cost $cx:=\sum(c(e)x(e)\colon e\in E)$.
(Since all stable assignments $x$ have the same size $|x|$, the function $c$ can be considered up to adding a constant; in particular, one may assume that $c$ is positive. Replacing $c$ by $-c$ leads to the problem of maximizing $cx$.)
Starting with the pioneering work [13], a well-known method of minimizing a linear function over the set of stable objects (in an appropriate stability model) consists in reduction to a minimization problem over the set of ideals or closed functions in the associated poset of rotations (assuming that such a poset exists and can be efficiently constructed). The latter problem is further reduced, by use of Picard’s method in [19], to the classical minimum cut problem. To make our description more self-contained, we now explain how this approach works in our case.
We first compute the cost of each rotation $\rho\in{\cal R}$, to be $c\rho:=\sum(c(e)\rho(e)\colon e\in E)$. For each $x\in{\cal S}$ and the corresponding closed function $\lambda=\omega(x)$, we have
$cx=cx^{\rm min}+\sum(\lambda(\rho)c\rho\colon\rho\in{\cal R}).$
Since problem ((7.6)) has an optimal solution $x\in{\cal S}$ for which $\omega(x)$ is fully closed, by ((7.5)), we come to the following minimization problem over the set of ideals in the finite poset $({\cal R},\lessdot)$, with the weight function $\zeta$ defined by $\zeta(\rho):=\tau(\rho)c\rho$ for $\rho\in{\cal R}$:
  1. (7.7)
    find an ideal $I\subset{\cal R}$ minimizing the total weight $\zeta(I):=\sum(\zeta(\rho)\colon\rho\in I)$.
Admitting a slightly more general setting (problem ()($\ast$)( ∗ )), we can consider a finite directed graph $H=(V_{H},E_{H})$ and a function $\zeta:V_{H}\to{\mathbb{R}}$ and ask for finding a closed set of vertices $X\subseteq V_{H}$ minimizing $\zeta(X):=\sum(\zeta(v)\colon v\in X)$ (where $X$ is called closed if there is no edge going from $V_{H}-X$ to $X$).
Following [19], problem ()($\ast$)( ∗ ) is reduced to an instance of minimum cut problem on the directed graph $\widehat{H}=(\widehat{V},\widehat{E})$ and edge capacity function $h$ obtained from $H,\zeta$ by:
(a) adding two vertices, “source” $s$ and “sink” $t$;
(b) adding the set $E^{+}$ of edges $(s,v)$ for all vertices $v$ in $V^{+}:=\{v\in V_{H}\colon\zeta(v)>0\}$;
(c) adding the set $E^{-}$ of edges $(u,t)$ for all vertices $u$ in $V^{-}:=\{u\in V_{H}\colon\zeta(u)<0\}$;
(d) assigning $h(s,v):=\zeta(v)$ for all $v\in V^{+}$,   $h(u,t):=|\zeta(u)|$ for all $u\in V^{-}$, and $h(e):=\infty$ for all $e\in E_{H}$.
An $s$$t$ cut in $H$ is meant to be the set $\delta(A)$ of edges going from a set $A\subset\widehat{V}$ such that $s\in A\not\ni t$ to its complement $\widehat{V}-A$, and its capacity is $h(\delta(A)):=\sum(h(e)\colon e\in\delta(A))$. One shows that $\delta(A)$ is a minimum capacity cut among all $s$$t$ cuts if and only if $X:=V_{H}-A$ is a closed set in $H$ with $\zeta(X)$ minimum.
Indeed, for an $s$$t$ cut $\delta(A)$, the value $h(\delta(A))$ is finite if and only if $\delta(A)$ contains no edges from $H$. The latter implies that $\delta(A)\subseteq E^{+}\cup E^{-}$, and $X$ is closed. Then
$h(\delta(A))=h(\delta(A)\cap E^{+})+h(\delta(A)\cap E^{-})\\ =\zeta(X\cap V^{+})+\sum(|\zeta(u)|\colon u\in(V_{H}-X)\cap V^{-})\\ =\zeta(X\cap V^{+})+\zeta(X\cap V^{-})-\zeta(V^{-})=\zeta(X)-\zeta(V^{-}).$
So $\zeta(X)$ differs from $h(\delta(A))$ by the constant $\zeta(V^{-})$, and the desired property follows. Summing up the above reasonings and constructions, we obtain
Theorem 7.1
The minimum cost stable assignment problem ((7.6)) is solvable in strongly polynomial time.
Remark 5. As is mentioned in the Introduction, the (strongly) polynomial-time solvability of ((7.6)) implies that the separation problem for ${\cal S}$ is solvable in (strongly) polynomial time as well, due to the polynomial-time equivalence between optimization and separation problems (in a wide class of such problems) established in [10, Theorems 6.4.9, 6.6.5]. (Recall that the separation problem for ${\cal S}$ is: given a vector $\xi\in{\mathbb{R}}^{E}$, decide whether $\xi$ is contained in ${\rm conv}({\cal S})$, and if not, find a hyperplane $H$ in ${\mathbb{R}}^{E}$ separating $\xi$ and ${\cal S}$, i.e. such that one component (half-space) in ${\mathbb{R}}^{E}-H$ contains $\xi$ but no element of ${\cal S}$.) Moreover, when $\xi\notin{\rm conv}({\cal S})$, one can find, in (strongly) polynomial time, a separating hyperplane including a facet of ${\rm conv}({\cal S})$ (cf. [10, Th. 6.5.16]).
In light of this, one may ask: (a) can ${\rm conv}({\cal S})$ be explicitly described via a system of linear constraints, or (b) can one characterize the facet structure of ${\rm conv}({\cal S})$ in “good combinatorial terms”? Both questions are affirmatively answered in the simplest case, the stable marriage problem, in [23]. An affirmative answer to (a) is given in [18] for the unconstrained SAP, and in [7] for the Boolean version of [1] in the case when at least one side of the market (viz. vertex part of $G$) is subject to quota restrictions and the choice functions are given via oracle calls. The latter generalizes known many-to-many stability problems. As to our SMP model, both questions are open at present.